OpenStudy (anonymous):

The supplement of an angle measures 40degrees more than twice its compliement find the measure of the angle? The compliement =90 -x The suppliment=180-x 40 degrees more than twice the compliement =2(90-x)+40=180-2x+40=220-2x so, 180-x=220-2x which would mean that the measure would still be 40 degrees is this correct?

OpenStudy (anonymous):

or 2x-x=220-180 which will still bring you back to a degree of 40

OpenStudy (cwrw238):

2(90-x) + 40 = 180- x 180 - 2x = 140 - x x = 40 right